# LLMs Can’t Handle Peer Pressure: Crumbling under Multi-Agent Social Interactions

### v2

@@ 2 Related Work
Cognitive Biases in Multi-Agent Systems
Recent studies show that AI systems, especially large language models (LLMs), can develop and even amplify human-like cognitive biases, affecting reasoning and decision-making in both individuals and groups (Chen et al. 2024a; Shaki, Kraus, and Wooldridge 2023). For example, agents align with group consensus even when it’s incorrect (Liu et al. 2025; Cho, Guntuku, and Ungar 2025). However, these studies do not address how to reduce such behaviours or manage complex social dynamics. To bridge this gap, Kairos offers a framework to evaluate how fine-tuning and reinforcement learning affect model performance in socially interactive settings.
Existing Benchmarks for Conformity
Existing benchmarks examine conformity bias in LLMs mainly through factual or logical QA and prompt-based debiasing (Zhu et al. 2025; Weng, Chen, and Wang 2025). While these methods measure alignment with ground truth, they neglect broader cognitive skills like creative problem-solving and social reasoning. Our Kairos benchmark fills this gap by extending evaluation to creativity and social understanding. It also provides fine-grained control over interactions, allowing systematic manipulation of social variables and deeper analysis of model behaviour in multi-agent settings.
We present a multi-agent interactive benchmark called Kairos designed to simulate socially grounded scenarios and assess LLM behaviour within them. Unlike benchmarks focused solely on conformity (Weng, Chen, and Wang 2025), Kairos targets how LLMs interpret, utilise, or resist signals from other agents, even when they are clearly unreliable, based on perceived reliability, current social context, and self-confidence. We begin by describing the details of collecting dynamic evaluation data in Section 3.1 and then explaining the corresponding metrics in Section 3.2.
3.1 Data Collection
A detailed distribution of the training and evaluation datasets is presented in Figure 1, covering various domains—Reasoning, Knowledge, Social, Common Sense, and Creativity—with careful partitioning to ensure no overlap and a clear distributional shift between training and evaluation phases. See more details in appendix A.
Figure 1: Left: Training dataset (N=10,000). Right: Test dataset (N=3,000). The inner ring groups tasks by category — Training: Reasoning 37.5%, Knowledge 21.1%, Social 20.8%, Creativity 20.5%; Test: Reasoning 33.3%, Knowledge 22.3%, Common Sense 22.2%, Creativity 22.2%. The outer ring breaks each category into individual datasets; wedge labels give original instance counts.
@@ 3.1 Data Collection
To evaluate models in realistic social dynamics, we collect datasets across four categories: reasoning, knowledge, common sense, and creativity.
Reasoning: Combines logic reasoning (Weng, Chen, and Wang 2025), filtered BIG-Bench Hard (Suzgun et al. 2022), code execution tasks from LiveCodeBench (Jain et al. 2024), and level 4–5 problems from MATH-500 (Lightman et al. 2023). Knowledge: Uses TruthfulQA (Lin, Hilton, and Evans 2021) and MMLU-Pro (Wang et al. 2024) for factual and broad-domain knowledge evaluation. Common Sense: Includes CommonsenseQA 2.0 (Talmor et al. 2022) and Social IQ (Sap et al. 2019) for intuitive and socially grounded reasoning. Creativity: Incorporates MacGyver (Tian et al. 2024) for situational problem-solving and BrainTeaser (Jiang et al. 2023) for lateral thinking.
All tasks are reformulated into multiple-choice question answering (MCQA) format. For open-ended datasets (e.g., LiveCodeBench, MATH-500), distractor options are generated via Llama3.1-8B and they were checked to ensure they are distinct from the ground truths. The MacGyver dataset is reframed as a binary judgment task to retain its creative challenge.
Dynamic Evaluation Dataset Construction.
To evaluate a model’s robustness in socially rich settings, we first identify its underlying beliefs: its preferred answer and the associated confidence for each benchmark question. This enables us to construct a tailored evaluation scenario that targets the model’s own epistemic commitments, rather than relying on a fixed benchmark. The resulting Kairos is therefore dynamically constructed for each model, adapting to its responses and beliefs to stress-test its social reasoning. The construction of Kairos proceeds in two main stages.
Step 1: Extracting the Model’s Original Beliefs
We begin by presenting the model with the original benchmark question and record its direct output, referred to as the model’s original answer, which serves as its stated belief. To additionally estimate the model’s confidence in this belief, we apply a Monte Carlo dropout-based uncertainty estimation procedure. Specifically, we sample $T$ stochastic forward passes to obtain predictive distributions:
$\left\{\mathbf{p}_{t}=p(y\mid\mathbf{x},\theta_{t})\right\}_{t=1}^{T}$
where $\mathbf{x}$ is the input, $y$ the predicted label, and $\theta\_t$ the model parameters under dropout at step $t$. The confidence is quantified via the predictive entropy of the averaged distribution $\bar{\mathbf{p}}$:
$\mathcal{H}[\bar{\mathbf{p}}]=-\sum_{k=1}^{K}\bar{p}_{k}\log\bar{p}_{k}$
Samples are categorized into high or low confidence by comparing their entropy to the dataset-wide median: those above the median are labeled low-confidence; those below, high-confidence.
Step 2: Simulating Social Scenarios with Targeted Interventions
Once the model’s belief, its predicted answer and associated confidence have been established, we construct a multi-agent simulation to examine how it responds to social influence. Each simulation consists of two components: simulated interaction history and current question round.
The interaction history mimics prior rounds of a quiz-style contest. For each past round, we provide the previous questions, the model’s own answers, and the responses of peer agents. This context allows us to simulate the buildup of agent-specific trust based on how consistently each peer has aligned with the evaluated model’s prior answers.
In the current round, the model is given a new question along with responses from peer agents that are intentionally crafted to either align with or challenge its previously expressed belief. This alignment is defined along three behavioural modes: support, oppose-hard, and oppose-easy. If the model’s original answer is correct, support agents reinforce it by repeating the correct answer, oppose-hard agents challenge it by selecting the most plausible incorrect option (i.e., the incorrect choice with the highest model-predicted probability), and oppose-easy agents offer minimal resistance by selecting the least plausible incorrect answer. Conversely, if the model’s answer is incorrect, support agents echo the same incorrect response, oppose-hard agents provide a different but highly plausible incorrect alternative, and oppose-easy agents present the correct answer. This targeted construction allows us to simulate varying degrees of social pressure and systematically assess how the model responds to both reinforcing and contradicting social signals.
To simulate diverse social scenarios, we uniformly vary the following hyperparameters to construct the Kairos benchmark:
  • Peer Agent Behaviour in the current round: Each agent either supports, opposes-hard, or opposes-easy the model’s belief identified in Step 1. This allows us to control the direction and strength of social pressure.
  • Number of Peer Agents: Varies the size of the social group influencing the model.
  • Peer Agent Trust Level: Ranges from $0\%,25\%,50\%,75\%,100\%$} and reflects each peer’s perceived reliability. Trust is earned when a peer’s past answers align with the model’s; it is diminished when they diverge. By varying the frequency of agreement across simulated history, we control the trust profile of each agent.
This setup enables a structured exploration of how factors like peer agreement, trustworthiness, and group size influence the model’s behaviour under social pressure.
@@
7:  Sample trust level $t_{q}\in\{0,25,50,75,100\}$
8:  Randomly choose trust rounds $\mathcal{R}_{\text{trust}}\subseteq\{1,\dots,R\}$ with $|\mathcal{R}_{\text{trust}}|=\frac{t_{q}}{100}\cdot R$
@@
10:   if $r\in\mathcal{R}_{\text{trust}}$ then
@@
3.2 Evaluation Metrics
Accuracy and Robustness.
Utility & Resistance.
[…]
Then, the utility of model $M$ is the fraction of originally incorrect samples that are corrected:
$U_{M}=\frac{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=0\wedge y_{i}=1\}}{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=0\}}\;\in[0,1].$ (2)
Similarly, the resistance of model $M$ is the fraction of originally correct samples that are maintained:
$R_{M}=\frac{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=1\wedge y_{i}=1\}}{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=1\}}\;\in[0,1].$ (3)
4 Experimental Setup
4.1 Mitigation Strategies
To improve the robustness of LLMs in socially interactive environments, we investigate three categories of mitigation strategies: Prompting, Supervised Fine-Tuning (SFT), and Reinforcement Learning using Group Relative Policy Optimisation (GRPO). Each approach aims to enhance the model’s ability to reason accurately while managing influence from other agents in multi-agent settings.
Prompting
Inspired by (Weng, Chen, and Wang 2025), we explore two different prompting strategies: Empowered and Reflective.
Supervised Fine-Tuning (SFT)
Reinforcement Learning
  1. 1.
    Multi-Agent vs. Non-Multi-Agent Context: In the MAS configuration, training inputs include the full history of prior questions and peer agent responses. In contrast, the Non-MAS setting removes all social context, allowing us to isolate the effects of agent interactions
  2. 2.
  3. 3.
  4. 4.
4.2 Training Dataset Construction for Mitigation Strategies
To avoid data contamination and ensure a clear train-test distributional shift, we build the training set from disjoint sources spanning four domains: reasoning, knowledge, social, and creativity.
Reasoning: Includes MathQA (Amini et al. 2019) for math word problems and Winogrande (Sakaguchi et al. 2021) for commonsense reasoning. Knowledge: Uses PopQA (Mallen et al. 2022) and SimpleQA (Wei et al. 2024) for factual question answering. Social: Combines Social (Yuan et al. 2024) and Moral Stories (Emelin et al. 2021) to promote understanding of social norms and moral reasoning. Creativity: Incorporates ProtoQA (Pro 2020) for prototypical reasoning and RiddleSense for metaphorical comprehension.
We process all datasets in the same way as during evaluation. For each interaction history, we ensure a balanced mix of round types (both support and oppose). In open-ended datasets such as SimpleQA (Wei et al. 2024), distractor answers are generated automatically. More details can be found in appendix A.
5 Results
Figure 3: Comparison of Original accuracy (solid bars) and Kairos accuracy (hatched bars) for 5 LLMs of different sizes and architectures under the different mitigation strategies. The robustness scores, $\text{O--K}~\Delta$ are presented over the bars.
5.1 Why Robustness Matters in Multi-Agent Settings
While accuracy improvements are valuable, robustness—the stability of performance between the Original and Kairos evaluations—is equally critical. A widening gap between these two metrics indicates that models are increasingly sensitive to social context. This fragility poses several challenges:
[…]
  • Vulnerability to Social Pressure: Our analysis shows that models typically lose more correct predictions than they gain from peer corrections, meaning they are more easily swayed into mistakes than helped out of them. Such behaviour resembles the human social dilemma and can propagate errors through the system.
  • Hidden Brittleness: Reinforcement learning often boosts surface-level accuracy but widens the Original–Kairos gap, implying overfitting to the evaluation protocol. Without robustness, accuracy gains do not translate into generalisation under distributional or social shifts.
  • System-level Risk: In collective settings, a socially induced error by a single agent can cascade through peers, amplifying misinformation and destabilising group results. This vulnerability can also be used for adversarial attacks on the system, causing safety issues.
Therefore, robustness should not be viewed as secondary to accuracy but as a core requirement for deploying LLMs in multi-agent environments. Improving accuracy without ensuring robustness risks building systems that perform well under controlled benchmarks yet fail under realistic, socially entangled conditions.
5.2 Main Results
GRPO Boosts Accuracy But Not Robustness.
As shown in Figure 3, training with GRPO consistently improves both original and Kairos performance over Supervised Fine-Tuning (SFT). On average, GRPO achieves a 12.3% increase in original accuracy and a 16.4% increase in Kairos accuracy compared to SFT. This confirms that reinforcement learning helps models better navigate socially interactive settings, where passive imitation is insufficient.
However, this gain comes with a trade-off in robustness, measured by $\text{O--K}~\Delta$. While GRPO models outperform in accuracy, they often exhibit increased performance drop from original to Kairos—especially when trained without Multi-Agent System (MAS) context—suggesting less robustness in generalization. This highlights a critical distinction: improved scores do not always imply better behavioral consistency under interaction-induced shifts.
MAS Context Enhances Both Performance and Robustness.
Prompting Improves Accuracy, But Harms Robustness in Most Models
Prompt-based strategies such as Empowered Prompting outperform SFT in both original and Kairos settings. Empowered Prompting achieves an average Kairos accuracy of 52.43%, compared to 48.16% for Reflective Prompting and 44.24% for SFT. Original accuracies follow a similar trend.
However, prompting methods tend to worsen robustness overall. On average, prompting models show lower $\text{O--K}~\Delta$ values, indicating greater sensitivity to shifts in evaluation protocols. Notably, Qwen2.5-7B is an exception, where Reflective Prompting improves robustness by +6%, suggesting model-dependent benefits. Conversely, LLaMA models exhibit a sharp drop in robustness (up to -20%) under Reflective Prompting, suggesting that added complexity may introduce harmful noise or instability depending on the underlying architecture.
NS-OR Configuration Yields Best Accuracy-Robustness Trade-off.
Filtering by Confidence Helps Accuracy but Harms Robustness.
Key Takeaways Strong performance requires balancing both accuracy and robustness, rather than optimizing for one alone. GRPO improves accuracy over SFT, but can reduce robustness in certain settings. Incorporating social context (MAS) during GRPO training boosts accuracy and, for larger models, enhances robustness, while smaller models may experience a drop. Debate-style reasoning enforced through system prompts or reward functions does not show improvements in robustness or accuracy, suggesting that models benefit more from simpler learning objectives.
6 Analysis
Our analysis spans two main dimensions: (1) the overall transition effect, capturing shifts in behaviour as models adapt within the MAS environment; and (2) the conditional influence of peer agent trust levels and actions in the current round. This framework highlights how different training strategies shape model sensitivity to social signals and structural dynamics in multi-agent coordination.
6.1 Overall Transition Effect
How LLMs Crumble under Multi‐Agent Social Interactions?
$R_{M}$ $=\Pr(\text{correct}\to\text{correct}),$
$1-R_{M}$ $=\Pr(\text{correct}\to\text{incorrect}),$
$U_{M}$ $=\Pr(\text{incorrect}\to\text{correct}),$
$1-U_{M}$ $=\Pr(\text{incorrect}\to\text{incorrect}).$
Then the post-interaction accuracy is
$A^{\prime}_{M}=p_{c}R_{M}+p_{i}U_{M},$
so the net change in accuracy is
$\Delta_{M}=A^{\prime}_{M}-p_{c}=p_{i}U_{M}-p_{c}(1-R_{M}).$
Since empirically $1-R_{M}>U_{M}$ across models and domains, and because $p_{c}$ typically exceeds $p_{i}$, we observe $\Delta_{M}<0$. In other words, the loss of correct predictions ($p_{c}(1-R_{M})$) outweighs the gains from correcting errors ($p_{i}U_{M}$), causing overall performance degradation under multi‐agent social interactions. See appendix C for more details.
[…]
6.2 Interaction Between Prior Trust and Social Signals
We observe systematic interactions between the prior trust level (0–100) and the current peer behaviour (support, oppose-hard, oppose-easy) across all tested model sizes, families, and training methods, as illustrated in Appendix 3. Increasing the trust level consistently strengthens models’ resistance in support conditions, reaching the highest resistance when trust is at 100. Conversely, resistance declines significantly in the oppose conditions as trust level grows, reaching its lowest point at trust level 100. For example, the original Qwen-14B-Instruct model’s resistance under support increases notably from 95.4 at trust level 0 to 99.2 at trust level 100, whereas under oppose-hard, resistance drops from 67.5 to 54.0, and under oppose-easy, from 73.6 to 51.2. Utility transitions show the exact opposite pattern: higher trust levels result in increased utility under the oppose conditions (e.g., Qwen-14B utility rising from 67.1 at trust level 0 to 72.5 at trust level 100 for oppose-easy) but reduced utility under the support condition (falling from 9.2 to 2.9). Notably, these behavioural shifts occur independently of confidence, which remains relatively constant across trust levels, suggesting that trust predominantly influences action selection rather than model certainty.
Regardless of trust level, the external signal type alone strongly affects model behavior. The support action consistently results in the highest resistance and lowest utility across all models, indicating a persistent, overly trusting stance toward supportive external information. Conversely, oppose-hard consistently yields the lowest performance in both resistance and utility, demonstrating that models struggle significantly when facing indirect or subtle corrective signals. Specifically, resistance under support is on average 23.4 counts higher than under oppose-easy and 31.7 counts higher than under oppose-hard, whereas utility shows the inverse pattern, reflecting an ingrained bias to favour supportive rather than critical external input.
Different training methods distinctly mediate these dynamics. GRPO improves both resistance and utility, particularly in challenging scenarios like oppose-hard, albeit at the expense of lower confidence. For instance, in Qwen-14B, GRPO raises resistance from 54.0 (Base) to 57.7 and utility from 13.9 to 16.7 under the oppose-hard condition, accompanied by a confidence drop of approximately 0.20. In contrast, SFT predominantly enhances performance in the simplest scenarios: significantly increasing resistance under support (e.g., Qwen-7B improving from 81.0 to 100.0) and utility under oppose-easy conditions (e.g., Llama-3B from 44.1 to 94.5). However, these improvements come at the cost of reduced effectiveness in genuinely challenging situations (oppose-hard), where performance can regress below the original instructive baseline. Overall, while models generally become more entrenched in existing beliefs with increased prior trust and struggle with subtle corrective cues, GRPO enhances robustness broadly, whereas SFT improves performance selectively, sacrificing generalisation under difficult correction signals.
7 Conclusion
In this work, we expand the notion of social bias to encompass how large language models (LLMs) form trust, resist misinformation, and selectively integrate peer input in social contexts to improve task performance—abilities critical for collaboration in future multi-agent systems (MAS). To investigate this, we introduce Kairos that systematically considers peer trust levels, peer actions, and the model’s own confidence. We find that current LLMs still struggle to resist external misinformation and incorporate peer input to correct their errors. In addition, we explore multiple training strategies—including prompting, supervised fine-tuning (SFT), and reinforcement learning under various configurations. Our results show that reinforcement learning with a simple outcome-based reward and unconstrained reasoning achieves the highest absolute performance in MAS settings, but at the cost of reduced relative robustness.
Appendix
Appendix A Data Construction Details
To comprehensively assess and train our models under realistic social dynamics, we organise our data into two main parts. First, we compile an evaluation collection covering four key dimensions—reasoning, knowledge, common sense, and creativity—to probe model behaviour across logical, factual, intuitive, and creative tasks. Second, we assemble a distinct training set drawn from separate sources in the same four domains, ensuring no overlap and a clear distributional shift. The following paragraphs describe the sources and processing steps for each of these collections in detail.
Evaluation Dataset Source.
Inspired by the limitations of prior work (Weng, Chen, and Wang 2025), we aim to move beyond datasets focused solely on logical reasoning. Instead, we expand the evaluation to more diverse scenarios to assess model performance under more realistic social dynamics. Specifically, we collect datasets from four categories: reasoning, knowledge, common sense, and creativity.
For the reasoning category, we include logic reasoning tasks from (Weng, Chen, and Wang 2025), filtered samples from the BIG-Bench Hard dataset (Suzgun et al. 2022), code execution and test output prediction tasks from LiveCodeBench (Jain et al. 2024), as well as level 4–5 problems from the MATH-500 dataset (Lightman et al. 2023). This selection ensures that model reasoning is not limited to logic inference, but also includes prediction and parallel mathematical thinking. In the knowledge category, we use TruthfulQA (Lin, Hilton, and Evans 2021), which focuses on fact-based questions and challenges the model’s parametric knowledge, and MMLU-Pro (Wang et al. 2024), which evaluates performance across a broad range of general-domain knowledge. To assess common sense, we employ CommonsenseQA 2.0 (Talmor et al. 2022) for basic factual commonsense, and additionally incorporate Social IQ (Sap et al. 2019) to simulate scenarios involving social behaviour, thereby broadening the coverage of everyday inference tasks. Finally, for the creativity dimension, we introduce datasets that evaluate the model’s ability to generate socially relevant and imaginative responses. We use MacGyver (Tian et al. 2024), which focuses on situational advice, and BrainTeaser (Jiang et al. 2023), which targets lateral thinking challenges. This multifaceted dataset collection enables a more comprehensive evaluation of models across logical, factual, intuitive, and imaginative aspects of social cognition.
We then convert each selected dataset into a multiple‐choice question answering (MCQA) format. For naturally multiple-choice datasets, this conversion is trivial. For open‐ended benchmarks such as LiveCodeBench (Jain et al. 2024) and MATH-500 (Lightman et al. 2023), we generate distractor options by prompting Llama3.1-8B with the original question and collecting distinct incorrect answers that differ from the ground truth. These automatically generated distractors are then subjected to both automated consistency checks and human review to ensure plausibility. For MacGyver (Tian et al. 2024), which is originally an open‐ended situational‐advice task, we recast it as a binary solvability judgment by asking the model to Judge whether the following problem is solvable. This reframing preserves the requirement for creative reasoning: a model capable of devising a valid solution should likewise assess its feasibility.
Training Dataset Construction.
To prevent data contamination and ensure a clear distributional shift between training and evaluation, we construct the training dataset from entirely distinct sources, categorized into four domains: reasoning, knowledge, social, and creativity.
In the reasoning domain, although part of the BBH dataset (Suzgun et al. 2022) overlaps with (Weng, Chen, and Wang 2025), we further incorporate MathQA (Amini et al. 2019), a dataset of multi-choice math word problems paired with executable programs, to strengthen the model’s mathematical reasoning capabilities, as well as Winogrande (Sakaguchi et al. 2021), a large-scale commonsense reasoning dataset based on Winograd-style pronoun resolution tasks. For the knowledge category, we utilise PopQA (Mallen et al. 2022), an open-domain QA benchmark grounded in Wikidata and designed to train factual recall, and SimpleQA (Wei et al. 2024), a short-form factual question answering dataset with single-entity answers. In the social domain, we use both Social (Yuan et al. 2024), which are datasets teaching reasoning about social norms and relationships, and Moral Stories (Emelin et al. 2021), which contains narratives requiring models to infer moral choices and socially acceptable actions. These datasets are selected to train reflective and socially sensitive reasoning behaviours. Finally, in the creativity domain, we use ProtoQA (Pro 2020), a dataset inspired by Family Feud where models must generate multiple plausible answers ranked by common sense likelihood, encouraging associative and prototypical reasoning. RiddleSense complements this by presenting riddles requiring indirect or metaphorical language understanding.
All datasets are processed analogously to the evaluation setup. Since SimpleQA (Wei et al. 2024) is an open-ended dataset without incorrect options, we generate distractor answers for each question using a pipeline similar to that used for LiveCodeBench (Jain et al. 2024) and Math-500 (Lightman et al. 2023) in the evaluation.
Appendix B Analyses

### v3

@@ 5 Related Work
5.1 Cognitive Biases in Multi-Agent Systems
Recent studies show that AI systems, especially large language models (LLMs), can develop and even amplify human-like cognitive biases, affecting reasoning and decision-making in both individuals and groups chen2024ai; shaki2023cognitive. For example, agents align with group consensus even when it’s incorrect liu2025exploringprosocialirrationalityllm; cho2025herd. However, these studies do not address how to reduce such behaviours or manage complex social dynamics. To bridge this gap, Kairos offers a framework to evaluate how fine-tuning and reinforcement learning affect model performance in socially interactive settings.
5.2 Existing Benchmarks for Conformity
Existing benchmarks examine conformity bias in LLMs mainly through factual or logical QA and prompt-based debiasing zhu2024conformity; wengwe. While these methods measure alignment with ground truth, they neglect broader cognitive skills like creative problem-solving and social reasoning. Our Kairos benchmark fills this gap by extending evaluation to creativity and social understanding. It also provides fine-grained control over interactions, allowing systematic manipulation of social variables and deeper analysis of model behaviour in multi-agent settings
6 Conclusion
In this work, we expand the notion of social bias to encompass how large language models (LLMs) form rapport, resist misinformation, and selectively integrate peer input in social contexts to improve task performance—abilities critical for collaboration in future multi-agent systems (MAS). To investigate this, we introduce Kairos that systematically considers peer rapport levels, peer actions, and the model’s own confidence. We find that current LLMs still struggle to resist external misinformation and incorporate peer input to correct their errors. In addition, we explore multiple training strategies—including prompting, supervised fine-tuning (SFT), and reinforcement learning under various configurations. Our results show that reinforcement learning with a simple outcome-based reward and unconstrained reasoning achieves the highest absolute performance in MAS settings, but at the cost of reduced relative robustness.
Appendix
Appendix A Data Construction Details
To comprehensively assess and train our models under realistic social dynamics, we organise our data into two main parts. First, we compile an evaluation collection covering four key dimensions—reasoning, knowledge, Social, and creativity—to probe model behaviour across logical, factual, intuitive, and creative tasks. Second, we assemble a distinct training set drawn from separate sources in the same four domains, ensuring no overlap and a clear distributional shift. The following paragraphs describe the sources and processing steps for each of these collections in detail.
@@ Appendix A Data Construction Details
Inspired by the limitations of prior work wengwe, we aim to move beyond datasets focused solely on logical reasoning. Instead, we expand the evaluation to more diverse scenarios to assess model performance under more realistic social dynamics. Specifically, we collect datasets from four categories: reasoning, knowledge, Social, and creativity.
For the reasoning category, we include logic reasoning tasks from wengwe, filtered samples from the BIG-Bench Hard dataset bbh, code execution and test output prediction tasks from LiveCodeBench jain2024livecodebench, as well as level 4–5 problems from the MATH-500 dataset lightman2023lets. This selection ensures that model reasoning is not limited to logic inference, but also includes prediction and parallel mathematical thinking. In the knowledge category, we use TruthfulQA lin2021truthfulqa, which focuses on fact-based questions and challenges the model’s parametric knowledge, and MMLU-Pro wang2024mmlu, which evaluates performance across a broad range of general-domain knowledge. To assess Social, we employ CommonsenseQA 2.0 talmor2022commonsenseqa for basic factual commonsense, and additionally incorporate Social IQ sap2019socialiqacommonsensereasoningsocial to simulate scenarios involving social behaviour, thereby broadening the coverage of everyday inference tasks. Finally, for the creativity dimension, we introduce datasets that evaluate the model’s ability to generate socially relevant and imaginative responses. We use MacGyver tian2023macgyver, which focuses on situational advice, and BrainTeaser jiang2023brainteaser, which targets lateral thinking challenges. This multifaceted dataset collection enables a more comprehensive evaluation of models across logical, factual, intuitive, and imaginative aspects of social cognition.
We then convert each selected dataset into a multiple‐choice question answering (MCQA) format. For naturally multiple-choice datasets, this conversion is trivial. For open‐ended benchmarks such as LiveCodeBench jain2024livecodebench and MATH-500 lightman2023lets, we generate distractor options by prompting Llama3.1-8B with the original question and collecting distinct incorrect answers that differ from the ground truth. These automatically generated distractors are then subjected to both automated consistency checks and human review to ensure plausibility. For MacGyver tian2023macgyver, which is originally an open‐ended situational‐advice task, we recast it as a binary solvability judgment by asking the model to Judge whether the following problem is solvable. This reframing preserves the requirement for creative reasoning: a model capable of devising a valid solution should likewise assess its feasibility.
Training Dataset Construction.
To prevent data contamination and ensure a clear distributional shift between training and evaluation, we construct the training dataset from distinct sources, categorised into four domains: reasoning, knowledge, social, and creativity.
In the reasoning domain, we continue to use a portion of source questions from BBH bbh and LiveCodeBench jain2024livecodebench, as both benchmarks offer representative and diverse reasoning challenges. Although they appear in both training and evaluation, we enforce a strict train–test split to avoid any data leakage. For BBH, the evaluation set contains a filtered subset of logical-reasoning tasks, whereas the training set uses a separate collection of reasoning samples. LiveCodeBench is handled analogously, with mutually exclusive problem sets ensuring that no prompt or solution overlaps between phases. Beyond these datasets, we also incorporate MathQA amini-etal-2019-mathqa, a multiple-choice math word problem dataset paired with executable programs, to strengthen mathematical reasoning, as well as Winogrande ai2winogrande, a large-scale commonsense reasoning benchmark based on Winograd-style pronoun resolution tasks. For the knowledge category, we utilise PopQA mallen2023llmmemorization, an open-domain QA benchmark grounded in Wikidata and designed to train factual recall, and SimpleQA wei2024measuringshortformfactualitylarge, a short-form factual question answering dataset with single-entity answers. In the social domain, we use both Social yuan2024measuring, which are datasets teaching reasoning about social norms and relationships, and Moral Stories Emelin2021MoralSS, which contains narratives requiring models to infer moral choices and socially acceptable actions. These datasets are selected to train reflective and socially sensitive reasoning behaviours. Finally, in the creativity domain, we use ProtoQA ProtoQA, a dataset inspired by Family Feud where models must generate multiple plausible answers ranked by common sense likelihood, encouraging associative and prototypical reasoning. RiddleSense complements this by presenting riddles requiring indirect or metaphorical language understanding.
All datasets are processed analogously to the evaluation setup. Since SimpleQA wei2024measuringshortformfactualitylarge is an open-ended dataset without incorrect options, we generate distractor answers for each question using a pipeline similar to that used for LiveCodeBench jain2024livecodebench and Math-500 lightman2023lets in the evaluation.
Dynamic Kairos Construction
To complement the description in Section 2.1, We outlines the dynamic data construction pipeline, showing how different configurations are varied and how the benchmark is tailored to target a model’s specific parametric beliefs.
@@
7:  Sample rapport level $t_{q}\in\{0,25,50,75,100\}$
8:  Randomly choose rapport rounds $\mathcal{R}_{\text{rapport}}\subseteq\{1,\dots,R\}$ with $|\mathcal{R}_{\text{rapport}}|=\frac{t_{q}}{100}\cdot R$
@@
10:   if $r\in\mathcal{R}_{\text{rapport}}$ then
@@
Appendix B Implementation details